# PATS: Proficiency-Aware Temporal Sampling for Multi-View Sports Skill Assessment

### v2

@@ II-A Temporal Sampling Methods
Temporal sampling strategies have evolved from uniform approaches to sophisticated content-aware methods. While Temporal Segment Networks (TSN) [2] achieves computational efficiency through sparse sampling, it sacrifices temporal continuity essential for skill assessment. Recent advances including motion-density methods, temporal correspondence techniques like sandwich sampling [3], channel sampling strategies [4], and temporal contextualization [5] have primarily focused on general action recognition, with limited consideration for the specialized requirements of proficiency assessment where temporal dynamics distinguish expert from novice performance.
@@ II-B Action Quality Assessment
Action Quality Assessment (AQA) evaluates how well actions are performed, requiring sensitivity to subtle performance differences [6]. Modern methods have evolved from handcrafted features to deep learning architectures using pre-trained backbones like I3D [7] and transformers [8]. Recent advances include spatial-aware modeling [9], temporal-aware approaches leveraging procedural action structure [10], and quality prediction strategies [11, 12].
Multi-modal approaches integrate skeleton features [13, 14] and audio information [15, 16] for enhanced understanding. Multi-view methods like SkillFormer [1] and EgoPulseFormer [17] fuse egocentric and exocentric features using cross-attention mechanisms and physiological signals, respectively. Datasets such as EgoExoLearn [18] and EgoExo4D [19] provide benchmarks for multi-perspective skill evaluation.
However, existing methods inherit temporal sampling limitations, failing to preserve the temporal dynamics that distinguish expert from novice performance. Our work addresses these limitations through specialized sampling strategies designed for skill assessment scenarios.
@@ III-A Proficiency-Aware Temporal Sampling (PATS)
We introduce Proficiency-Aware Temporal Sampling (PATS), a novel temporal sampling strategy designed to preserve the sequential nature of skilled performances for accurate proficiency assessment. Unlike traditional uniform sampling methods that randomly select frames across a video, PATS extracts continuous temporal segments that maintain the natural flow of athletic movements and performance dynamics.
@@ III-A2 Adaptive Segment Positioning
Given an input video of total duration $T$ seconds with $N_{total}$ frames captured at frame rate $fps$, PATS first determines the effective segment duration to prevent temporal overlap and ensure adequate spacing:
@@
The 0.8 scaling factor ensures sufficient buffer space between segments while maximizing temporal coverage.
@@ III-A3 Frame Allocation and Extraction
For each segment $i$ with temporal boundaries $[t_{start,i},t_{start,i}+d_{s,eff}]$, the corresponding frame indices are computed as:
@@ III-A4 Robust Edge Case Handling
The final output is a sorted list of exactly $N_{target}$ frame indices clipped to $[0,N_{total}-1]$, ensuring robust behavior across all input conditions.
@@ III-B Integration with SkillFormer
PATS is designed as a general video sampling strategy that can enhance any proficiency estimation architecture relying on temporal dynamics. We integrate PATS with SkillFormer [1] for three key reasons: (1) SkillFormer represents the current state-of-the-art for proficiency estimation on EgoExo4D to the best of our knowledge, (2) PATS specifically targets proficiency estimation tasks where temporal continuity distinguishes expert from novice performance, and (3) both approaches focus on sports domains, which constitute a significant portion of EgoExo4D scenarios.
PATS enhances SkillFormer’s multi-view architecture by preserving temporal continuity essential for effective cross-view fusion. SkillFormer’s CrossViewFusion module employs multi-head cross-attention mechanisms to correlate features across camera perspectives, requiring meaningful temporal relationships to capture performance nuances. By providing continuous temporal segments rather than sparse frames, PATS enables SkillFormer’s attention computations to effectively analyze movements across multiple views.
PATS seamlessly replaces SkillFormer’s uniform temporal sampling while maintaining full compatibility with the original architecture. The preserved temporal structure enhances proficiency assessment across diverse sporting activities without requiring architectural modifications.
@@ IV-A Dataset
Following established protocols [1, 17], we adopt the official dataset partitions, reserving 10% of the training set for validation and conducting final evaluation on the official held-out validation set. The dataset’s synchronized multi-view footage provides an ideal framework for assessing temporal sampling strategies in skill assessment scenarios.
@@ IV-B Implementation Details
To ensure fair comparison, we maintain identical training configurations to the original SkillFormer implementation [1]. All models are initialized from a TimeSformer backbone [20], pre-trained on Kinetics-600 [21], and fine-tuned for 4 epochs using AdamW optimizer with weight decay of 0.01 and LoRA adaptation. We adopt SkillFormer’s established hyperparameters for each experimental setup, including projector hidden dimensions, LoRA rank and scaling factors, learning rate, and batch size configurations.
@@ IV-C PATS Hyperparameter Selection Rationale
Duration configuration explores 1-second and 3-second segments based on sports biomechanics principles: short duration (1s) captures rapid execution phases in high-frequency activities, while standard duration (3s) encompasses complete movement sequences optimal for most scenarios.
@@
Exo only External observation scenarios (Cooking)
EgoExo Team sports & complex interactions (Basketball, Dancing)
@@ V-C Scenario-Specific Analysis
Table IV shows the optimal configurations identified for each scenario, revealing domain-specific preferences that align with skill characteristics. Basketball achieves the highest absolute accuracy (78.76%) using rapid multi-view sampling (5.33 FPS, 2 segments) that preserves continuous gameplay flow. Music reaches 74.14% through fine-grained egocentric capture (0.89 FPS, 12 segments) suited to sequential note execution patterns. Cooking performs optimally with exocentric-only views (60.53%) using high-frequency sampling (4.00 FPS, 8 segments) for technique visibility, while bouldering achieves 42.31% with rapid egocentric sampling (5.33 FPS, 2 segments) focusing on proprioceptive feedback.
Table V provides detailed per-scenario comparisons across all methods and viewing configurations. PATS delivers substantial improvements in several domains: bouldering shows the largest gain (+26.22% over SkillFormer), music improves by +2.39%, and basketball by +1.13%, with the latter reaching the best overall performance across all activities. These results demonstrate PATS’ effectiveness for proprioceptive skills requiring precise temporal coordination.
However, PATS shows mixed results in some scenarios. For dancing, while PATS improves over SkillFormer (26.50% vs 20.51%), it remains below baseline methods (55.65%), suggesting that this domain may require alternative PATS configurations with different parameter combinations beyond those explored in our grid search. For soccer, PATS shows a decline specifically in egocentric view accuracy, dropping from 70.83% with SkillFormer to 66.67% with PATS, while maintaining comparable performance in other viewing configurations, indicating that egocentric temporal sampling may be less suitable for soccer fundamental assessment.

### v3

@@ II-A Temporal Sampling Methods
Temporal sampling determines which frames to extract from video sequences for processing. Traditional approaches have evolved from uniform methods to sophisticated content-aware strategies. While Temporal Segment Networks (TSN) [2] achieve computational efficiency through sparse sampling (selecting few frames distributed across the video duration), they sacrifice temporal continuity essential for skill assessment. Recent advances including motion-density methods, temporal correspondence techniques like sandwich sampling [3], channel sampling strategies [4], and temporal contextualization [5] have primarily focused on general action recognition, with limited consideration for the specialized requirements of proficiency assessment where temporal dynamics distinguish expert from novice performance.
@@ II-B Action Quality Assessment
Action Quality Assessment (AQA) evaluates how well actions are performed, requiring sensitivity to subtle performance differences [6]. Modern methods have evolved from handcrafted features to deep learning architectures using pre-trained backbones like I3D [7] and transformers [8]. Recent advances include spatial-aware modeling [9], temporal-aware approaches [10], and quality prediction strategies [11, 12].
Multi-modal approaches integrate skeleton features [13, 14] and audio information [15, 16] for enhanced understanding. Multi-view methods show particular promise: SkillFormer [1] processes synchronized egocentric and exocentric videos through a TimeSformer backbone with LoRA fine-tuning, using CrossViewFusion with multi-head cross-attention for view-specific feature fusion. However, it relies on uniform sampling, which fails to preserve complete temporal dynamics. EgoPulseFormer [17] leverages physiological signals from eye-tracking cameras. Datasets such as EgoExoLearn [18] and EgoExo4D [19] provide multi-perspective benchmarks.
Existing methods inherit temporal sampling limitations, failing to preserve the temporal dynamics that distinguish expert from novice performance. Our work addresses these limitations through specialized sampling strategies designed for skill assessment scenarios.
Figure 1: In this configuration, PATS extracts $N_{target}=32$ frames from $N_{s}=2$ continuous temporal segments of duration $d_{s}=3s$ from a 10 s video. Within each segment, $\lfloor N_{target}/N_{s}\rfloor=16$ frames are sampled uniformly (red vertical lines), preserving temporal continuity within segments. Segment positioning with automatic spacing prevents overlap and ensure comprehensive temporal coverage. This configuration is used in the basketball and bouldering domains.
@@ III-A Proficiency-Aware Temporal Sampling (PATS)
We introduce Proficiency-Aware Temporal Sampling (PATS), a novel temporal sampling strategy designed to preserve the sequential nature of skilled performances for accurate proficiency assessment. Unlike traditional uniform sampling methods that randomly select frames across a video, PATS extracts continuous temporal segments that maintain the natural flow of athletic movements and performance dynamics (1).
@@ III-A2 Adaptive Segment Positioning
Given an input video of total duration $T$ seconds with $N_{total}$ frames captured at frame rate fpsfps, PATS first determines the effective segment duration to prevent temporal overlap and ensure adequate spacing:
@@
The 0.8 scaling factor limits each segment to 80% of its theoretical maximum duration, creating natural spacing between segments and preventing temporal overlap.
@@ III-A3 Frame Allocation and Extraction
For each segment $i$ with temporal boundaries $t_{start,i},t_{start,i}+d_{s,eff}]$, the corresponding frame indices are computed as:
@@ III-A4 Robust Edge Case Handling
The final output is a sorted list of exactly $N_{target}$ frame indices clipped to $0,N_{total}-1]$, ensuring robust behavior across all input conditions.
@@ III-B Integration with SkillFormer
We integrate PATS with SkillFormer [1] as it represents the current state-of-the-art for proficiency estimation on EgoExo4D, with both approaches targeting sports domains where temporal continuity distinguishes expert from novice performance.
PATS enhances SkillFormer’s multi-view architecture by preserving temporal continuity essential for cross-view fusion. SkillFormer’s CrossViewFusion module requires meaningful temporal relationships to capture performance nuances across camera perspectives. By providing continuous temporal segments rather than sparse frames, PATS enables more effective attention computations for multi-view movement analysis.
PATS seamlessly replaces SkillFormer’s original sampling while maintaining full architectural compatibility, enhancing proficiency assessment without requiring modifications.
@@ IV-A Dataset
Following established protocols [1, 17], we adopt the official dataset partitions, reserving 10% of the training set for validation and conducting final evaluation on the official held-out validation set.
@@ IV-B Implementation Details
To ensure fair comparison, we maintain identical training configurations to the original SkillFormer implementation [1]. All models are initialized from a TimeSformer backbone [20], pre-trained on Kinetics-600 [21], and fine-tuned for 4 epochs using AdamW optimizer with weight decay of 0.01 and LoRA adaptation. We adopt SkillFormer’s hyperparameters for each experimental setup.
@@ IV-C PATS Hyperparameter Selection Rationale
Duration configuration explores 1 s and 3 s segments based on sports biomechanics principles: short duration (1 s) captures rapid execution phases in high-frequency activities, while standard duration (3 s) encompasses complete movement sequences optimal for most scenarios.
@@
Exos only External observation scenarios (Cooking)
Ego+Exos Team sports & complex interactions (Basketball, Dancing)
@@ V-C Scenario-Specific Analysis
Table IV shows optimal configurations for each scenario, revealing domain-specific preferences. Basketball achieves the highest accuracy (78.76%) using rapid multi-view sampling (5.33 FPS, 2 segments), while music reaches 74.14% through fine-grained egocentric capture (0.89 FPS, 12 segments). Cooking performs best with exocentric-only views (60.53%) using high-frequency sampling, while bouldering achieves 42.31% with rapid egocentric sampling.
Table V provides detailed per-scenario comparisons. PATS delivers substantial improvements in several domains: bouldering shows the largest gain (+26.22% over SkillFormer), music improves by +2.39%, and basketball by +1.13%. These results demonstrate PATS’ effectiveness for proprioceptive skills requiring precise temporal coordination.
However, PATS shows mixed results in some scenarios. For dancing, while PATS improves over SkillFormer (26.50% vs 20.51%), it remains below baseline methods (55.65%), suggesting this domain may require alternative parameter combinations. For soccer, PATS shows a decline in egocentric view accuracy (66.67% vs 70.83%), indicating egocentric temporal sampling may be less suitable for soccer assessment.